# Provable Benefit of Annealed Langevin Monte Carlo for Non-log-concave Sampling

### v1

@@ 2.3 Wasserstein Distance and Curves of Probability Measures
We briefly introduce several fundamental concepts in optimal transport, and direct readers to authoritative textbooks [villani2008optimal, villani2021topics, ambrosio2008gradient, ambrosio2021lectures] for an in-depth exploration.
For two probability measures $\mu,\nu$ on $\mathbb{R}^{d}$ with finite second-order moments, the Wasserstein-2 (W2) distance between $\mu$ and $\nu$ is defined as
$W_{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\left(\int\left\|x-y\right\|^{2}\gamma({\rm d}x,{\rm d}y)\right)^{\frac{1}{2}},$
@@
$\left|\dot{\rho}\right|_{t}:=\lim_{\delta\to 0}\frac{W_{2}(\rho_{t+\delta},\rho_{t})}{\left|\delta\right|},$
which can be interpreted as the “speed” of this curve. If $\left|\dot{\rho}\right|_{t}$ exists and is finite for all $t\in[a,b]$, we say that $\rho$ is absolutely continuous (AC). The metric derivative and the continuity equation are closely related by the following important fact (see [ambrosio2008gradient, Theorem 8.3.1]):
@@
For an AC curve of probability measures $(\rho_{t})_{t\in[a,b]}$, any vector field $(v_{t})_{t\in[a,b]}$ that generates $(\rho_{t})_{t\in[a,b]}$ satisfies $\left|\dot{\rho}\right|_{t}\leq\left\|v_{t}\right\|_{L^{2}(\rho_{t})}$ for all $t\in[a,b]$. Moreover, the equality is reachable.
Finally, define the action of an AC curve of probability measures $(\rho_{t})_{t\in[a,b]}$ as ${\cal A}(\rho):=\int_{a}^{b}\left|\dot{\rho}\right|_{t}^{2}{\rm d}t$.
2.4 Isoperimetric Inequalities
A probability measure $\pi$ on $\mathbb{R}^{d}$ satisfies a log-Sobolev inequality (LSI) with constant $C$, or $C$-LSI, if for all $f\in C^{1}(\mathbb{R}^{d})$ with $\operatorname{\mathbb{E}}_{\pi}\left[f^{2}\right]>0$,
$\operatorname{\mathbb{E}}_{\pi}\left[f^{2}\log\frac{f^{2}}{\operatorname{\mathbb{E}}_{\pi}\left[f^{2}\right]}\right]\leq 2C\operatorname{\mathbb{E}}_{\pi}\left[\left\|\nabla f\right\|^{2}\right].$
It is worth noting that $\alpha$-strongly-log-concave distributions satisfy $\frac{1}{\alpha}$-LSI (Bakry-Émery theorem, [bakry2014analysis]). When $\pi\propto{\rm e}^{-V}$ satisfies $C$-LSI and the potential $V$ is $\beta$-smooth, it is established in [vempala2019rapid] that the LD converges exponentially fast in KL divergence, while the LMC also converges exponentially with a bias that vanishes when the step size approaches 00.
@@ 3 Problem Setup
In this paper, we consider the following mild assumption on the target distribution $\pi\propto{\rm e}^{-V}$ on $\mathbb{R}^{d}$, which is widely used in the field of sampling (see, e.g., [durmus2019analysis, vempala2019rapid, balasubramanian2022towards, chewi2022analysis]):
@@
The potential $V$ is $\beta$-smooth, and there exists a global minimizer $x_{*}$ of $V$ such that $\left\|x_{*}\right\|\leq R$. Moreover, $\pi$ has finite second-order moment.
Recall that the rationale behind annealing involves a gradual transition from a distribution that is easy to sample from to the more complex target distribution, which is crucial to our algorithm. Throughout this paper, we define the following curve of probability measures on $\mathbb{R}^{d}$:
$\pi_{\theta}\propto\exp\left(-\eta(\theta)V-\frac{\lambda(\theta)}{2}\left\|\cdot\right\|^{2}\right),~{}\theta\in[0,1].$ (3)
We use the term annealing schedule to refer to the functions $\eta(\cdot)$ and $\lambda(\cdot)$. They need to be differentiable and monotone, satisfying the boundary conditions $\eta_{0}=\eta(0)\nearrow\eta(1)=1$ and $\lambda_{0}=\lambda(0)\searrow\lambda(1)=0$. The values of $\eta_{0}\in[0,1]$ and $\lambda_{0}\in(1,{+\infty})$ will be determined later. This flexible interpolation scheme includes many general cases. For example, [brosse2018normalizing] and [ge2020estimating] used the schedule $\eta(\cdot)\equiv 1$, while [neal2001annealed] used the schedule $\lambda(\theta)=c(1-\eta(\theta))$. The key motivation for this interpolation is that when $\theta=0$, the quadratic term predominates, making the potential of $\pi_{0}$ strongly-convex with a moderate condition number, thus $\pi_{0}$ is easy to sample from; on the other hand, when $\theta=1$, $\pi_{1}$ is just the target distribution $\pi$. As $\theta$ gradually increases from 00 to $1$, the readily sampleable distribution $\pi_{0}$ slowly transform into the target distribution $\pi_{1}$.
In Lemma 5, we prove that $\pi_{\theta}$ also has finite second-order moment, so the W2 distance between $\pi_{\theta}$’s makes sense. We further make a mild assumption on the absolute continuity of the curve:
@@
The curve $(\pi_{\theta})_{\theta\in[0,1]}$ is AC with finite action ${\cal A}_{\eta,\lambda}(\pi):=\int_{0}^{1}\left|\dot{\pi}\right|_{\theta}^{2}{\rm d}\theta$.
To facilitate easy sampling from $\pi_{0}$, we consider two choices of the parameters $\eta_{0}$ and $\lambda_{0}$, which have the following complexity guarantees.
Lemma 3.
When $\eta_{0}=0$, $\pi_{0}=\operatorname{{\cal N}}\left(0,\frac{1}{\lambda_{0}}I\right)$ can be sampled directly; when $\eta_{0}\in(0,1]$, we choose $\lambda_{0}=\eta_{0}d\beta$, so that under Assumption 1, it takes $O\left(1\vee\log\frac{\eta_{0}\beta R^{2}}{d^{2}}\right)=\widetilde{O}\left(1\right)$ queries to the oracle of $V$ and $\nabla V$ in expectation to precisely sample from $\pi_{0}$ via rejection sampling.
See Appendix A for the rejection sampling algorithm as well as the full proof of this lemma. The parameter $R$ reflects prior knowledge about global minimizer(s) of the potential function $V$. Unless it is exceptionally large, indicating scarce prior information about the global minimizer(s) of $V$, this $\widetilde{O}\left(1\right)$ complexity is negligible compared to the overall complexity of sampling. In particular, when the exact location of a global minimizer of $V$ is known, we can adjust the potential $V$ so that 00 becomes a global minimizer, thereby reducing the complexity to $O(1)$.
Equipped with this foundational setup, we now proceed to introduce the annealed LD and annealed LMC algorithms.
@@ 4 Analysis of Annealed Langevin Diffusion
To elucidate the concept of annealing more clearly, we first consider the annealed Langevin diffusion (ALD) algorithm, which samples from the $\pi\propto{\rm e}^{-V}$ by running LD with a dynamically changing target distribution. For the purpose of this discussion, let us assume that we can exactly simulate any SDE with known drift and diffusion terms.
Fix an annealing schedule $\eta(\cdot),\lambda(\cdot)$ and choose a sufficiently long time $T$. We define a reparametrized curve of probability measures $\left(\widetilde{\pi}_{t}:=\pi_{t/T}\right)_{t\in[0,T]}$. Starting with an initial sample $X_{0}\sim\pi_{0}=\widetilde{\pi}_{0}$, we run the SDE
${\rm d}X_{t}=\nabla\log\widetilde{\pi}_{t}(X_{t}){\rm d}t+\sqrt{2}{\rm d}B_{t},~{}t\in[0,T],$ (4)
@@
When choosing $T=\frac{{\cal A}_{\eta,\lambda}(\pi)}{4\varepsilon^{2}}$, it follows that $\operatorname{KL}\left(\pi\middle\|\nu^{\rm ALD}\right)\leq\varepsilon^{2}$.
@@
Let $\mathbb{Q}$ be the path measure of ALD (Equation 4) initialized at $X_{0}\sim\widetilde{\pi}_{0}$, and define $\mathbb{P}$ as the path measure corresponding to the following reference SDE:
${\rm d}X_{t}=(\nabla\log\widetilde{\pi}_{t}+v_{t})(X_{t}){\rm d}t+\sqrt{2}{\rm d}B_{t},~{}X_{0}\sim\widetilde{\pi}_{0},~{}t\in[0,T].$ (5)
The vector field $v=(v_{t})_{t\in[0,T]}$ is designed such that $X_{t}\sim\widetilde{\pi}_{t}$ for all $t\in[0,T]$. According to the Fokker-Planck equation, $v$ must satisfy the PDE
@@
which means that $v$ generates $\left(\widetilde{\pi}_{t}:=\pi_{t/T}\right)_{t\in[0,T]}$. We can compute $\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)$ using Lemma 1:
$\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)=\frac{1}{4}\operatorname{\mathbb{E}}_{\mathbb{P}}\left[\int_{0}^{T}\left\|v_{t}(X_{t})\right\|^{2}{\rm d}t\right]=\frac{1}{4}\int_{0}^{T}\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}^{2}{\rm d}t.$
Leveraging Lemma 2, among all vector fields $v$ that generate $\left(\widetilde{\pi}_{t}:=\pi_{t/T}\right)_{t\in[0,T]}$, we can choose the one that minimizes $\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}$, thereby making $\left\|v_{t}\right\|_{L^{2}(\widetilde{\pi}_{t})}=\left|\dot{\widetilde{\pi}}\right|_{t}$, the metric derivative. With the reparameterization $\widetilde{\pi}_{t}=\pi_{t/T}$, we have the following relation by chain rule:
$\left|\dot{\widetilde{\pi}}\right|_{t}=\lim_{\delta\to 0}\frac{W_{2}(\widetilde{\pi}_{t+\delta},\widetilde{\pi}_{t})}{\left|\delta\right|}=\lim_{\delta\to 0}\frac{W_{2}(\pi_{(t+\delta)/T},\pi_{t/T})}{T\left|\delta/T\right|}=\frac{1}{T}\left|\dot{\pi}\right|_{t/T}.$
@@
$\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)=\frac{1}{4}\int_{0}^{T}\left|\dot{\widetilde{\pi}}\right|_{t}^{2}{\rm d}t=\frac{1}{4T}\int_{0}^{1}\left|\dot{\pi}\right|_{\theta}^{2}{\rm d}\theta=\frac{{\cal A}_{\eta,\lambda}(\pi)}{4T}.$
Finally, using the data-processing inequality ([chewi2022log, Theorem 1.5.3]), with $T=\frac{{\cal A}_{\eta,\lambda}(\pi)}{4\varepsilon^{2}}$,
$\operatorname{KL}\left(\pi\middle\|\nu^{\rm ALD}\right)=\operatorname{KL}\left(\mathbb{P}_{T}\middle\|\mathbb{Q}_{T}\right)\leq\operatorname{KL}\left(\mathbb{P}\middle\|\mathbb{Q}\right)=\varepsilon^{2},$
@@
Although we adopt a specific parametrization of the interpolating distribution (Equation 3), which is widely used in practice, the results in Theorem 1 are applicable to any curve of probability measures $(\pi_{\theta})_{\theta\in[0,1]}$ that bridge $\pi_{0}$ and $\pi_{1}=\pi$. This applicability extends as long as the closed forms of the drift and diffusion terms in the SDE, which precisely follow the path $(\pi_{\theta})_{\theta\in[0,1]}$, are known.
Let us delve deeper into the mechanics of the ALD. Although at time $t$ the SDE (Equation 4) targets the distribution $\widetilde{\pi}_{t}$, the distribution of $X_{t}$ does not precisely match $\widetilde{\pi}_{t}$. Nevertheless, by choosing a sufficiently long time $T$, we actually move on the curve $(\pi_{\theta})_{\theta\in[0,1]}$ sufficiently slowly, thus minimizing the discrepancy between the path measure of $(X_{t})_{t\in[0,T]}$ and the reference curve $(\widetilde{\pi}_{t})_{t\in[0,T]}$. By applying data-processing inequality, we can upper bound the error between the marginal distributions at time $T$ by the joint distributions of the two paths. In essence, moving more slowly, sampling more precisely.
Our analysis predominantly addresses the global error across the entire curve of probability measures, rather than focusing solely on the local error at time $T$. This approach is inspired by [dalalyan2012sparse] (see also [chewi2022log, Section 4.4]) and [chen2023sampling], and stands in contrast to the traditional isoperimetry-based analyses of LD (e.g., [vempala2019rapid, chewi2022analysis, balasubramanian2022towards]), which emphasize the decay of the KL divergence from the distribution of $X_{t}$ to the target distribution and require LSI to bound the time derivative of this quantity. Notably, the total time $T$ needed to run the SDE depends solely on the action of the curve $(\pi_{\theta})_{\theta\in[0,1]}$, obviating the need for assumptions related to log-concavity or isoperimetry.
We also remark that the ALD plays a significant role in another important subject known as non-equilibrium stochastic thermodynamics [seifert2012stochastic]. Recently, a refinement of the fluctuation theorem was discovered in [chen2020stochastic, fu2021maximal], stating that the irreversible entropy production in a stochastic thermodynamic system is equal to the ratio of a similar action integral as in Theorem 1 and the duration of the process $T$, resembling Theorem 1.

### v2

@@ 2.4 Wasserstein Distance and Curves of Probability Measures
We briefly introduce several fundamental concepts in optimal transport, and direct readers to authoritative textbooks (Villani, 2008; 2021; Ambrosio et al., 2008; 2021) for an in-depth exploration.
For two probability measures $\mu,\nu$ on $\mathbb{R}^{d}$ with finite second-order moments, the Wasserstein-2 ($\text{W}_{\text{2}}$) distance between $\mu$ and $\nu$ is defined as
$W_{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\left(\int\|x-y\|^{2}\gamma(\mathrm{d}x,\mathrm{d}y)\right)^{\frac{1}{2}},$
@@
$|\dot{\rho}|_{t}:=\lim_{\delta\to 0}\frac{W_{2}(\rho_{t+\delta},\rho_{t})}{|\delta|},$
which can be interpreted as the “speed” of this curve. If $|\dot{\rho}|_{t}$ exists and is finite for all $t\in[a,b]$, we say that $\rho$ is absolutely continuous (AC). AC is a fairly weak regularity condition on the curve of probability measures. In Lem. 4, we present a sufficient condition for AC as a formal statement.
The metric derivative and the continuity equation are closely related by the following important fact from Ambrosio et al. (2008, Theorems 8.3.1 and 8.4.5):
@@
For an AC curve of probability measures $(\rho_{t})_{t\in[a,b]}$, any vector field $(v_{t})_{t\in[a,b]}$ that generates $(\rho_{t})_{t\in[a,b]}$ satisfies $|\dot{\rho}|_{t}\leq\|v_{t}\|_{L^{2}(\rho_{t})}$ for a.e. $t\in[a,b]$. Moreover, there exists a unique vector field $(v^{*}_{t})_{t\in[a,b]}$ generating $(\rho_{t})_{t\in[a,b]}$ that satisfies $|\dot{\rho}|_{t}=\|v^{*}_{t}\|_{L^{2}(\rho_{t})}$ for a.e. $t\in[a,b]$.
Finally, we define the action of an AC curve of probability measures $(\rho_{t})_{t\in[a,b]}$ as $\int_{a}^{b}|\dot{\rho}|_{t}^{2}\mathrm{d}t$. As will be shown in the next section, the action is a key property characterizing the effectiveness of a curve in annealed sampling. The following lemma provides additional intuition about the action and may be helpful for interested readers.
Lemma 3.
Given an AC curve of probability measures $(\rho_{t})_{t\in[0,1]}$, and let ${\cal A}$ be its action. Then
  1. 1.
    ${\cal A}\geq W_{2}^{2}(\rho_{0},\rho_{1})$, and the equality is attained when $(\rho_{t})_{t\in[0,1]}$ is a constant-speed Wasserstein geodesic, i.e., let two random variables $(X_{0},X_{1})$ follow the optimal coupling of $(\rho_{0},\rho_{1})$ such that $\mathbb{E}\|X_{0}-X_{1}\|^{2}=W_{2}^{2}(\rho_{0},\rho_{1})$, and define $\rho_{t}$ as the law of $(1-t)X_{0}+tX_{1}$.
  2. 2.
    If $\rho_{t}$ satisfies $C_{\rm LSI}(\rho_{t})$-LSI for all $t$, then ${\cal A}\leq\int_{0}^{1}C_{\rm LSI}(\rho_{t})^{2}\|\partial_{t}\nabla\log\rho_{t}\|^{2}_{L^{2}(\rho_{t})}{\rm d}t.$
  3. 3.
    If $\rho_{t}$ satisfies $C_{\rm PI}(\rho_{t})$-PI for all $t$, then ${\cal A}\leq\int_{0}^{1}2C_{\rm PI}(\rho_{t})\|\partial_{t}\log\rho_{t}\|_{L^{2}(\rho_{t})}^{2}{\rm d}t$.
The proof of Lem. 3 can be found in App. A. The lower bound above can be derived by definition or via variational representations of action and ${\rm W}_{2}$ distance, while the two upper bounds relate the action to the isoperimetric inequalities, indicating that finite action is a weaker assumption than requiring the LSI or PI constants along the trajectory. In fact, these bounds may not be tight, as demonstrated by the following surprising example, in which the action is polynomial with respect to certain problem parameters and even independent of some of them, whereas the LSI and PI constants exhibit exponential dependence. Its proof is detailed in App. A.
Example 1.
Consider the curve that bridges a target distribution $\rho_{0}$ to a readily sampleable distribution $\rho_{1}$, in the form $\rho_{t}=\rho_{0}*\operatorname{{\cal N}}\left(0,2StI\right)$, $t\in[0,1]$, for some large positive number $S$. Let ${\cal A}$ be the action of this curve. Then
  1. 1.
    ${\cal A}=S\int_{0}^{S}\|\nabla\log p_{s}\|^{2}_{L^{2}(p_{s})}\mathrm{d}s$, where $p_{s}=\rho_{s/S}=\rho_{0}*\operatorname{{\cal N}}\left(0,2sI\right)$, $s\in[0,S]$.
  2. 2.
    In particular, when $\rho_{0}$ is a mixture of Gaussian distribution $\sum_{i=1}^{N}w_{i}\operatorname{{\cal N}}\left(\mu_{i},\sigma^{2}I\right)$, then ${\cal A}\leq\frac{Sd}{2}\log\left(1+\frac{2S}{\sigma^{2}}\right)$ is independent of the number of components $N$, the weights $w_{i}$, and the means $\mu_{i}$. However, in general, the LSI and PI constants may depend exponentially on $\max_{i,j}\|\mu_{i}-\mu_{j}\|$, the maximum distance between the means.
@@ 3 Problem Setup
Recall that the rationale behind annealing involves a gradual transition from $\pi_{0}$, a simple distribution that is easy to sample from, to $\pi_{1}=\pi$, the more complex target distribution. Throughout this paper, we define a curve of probability measures $\left(\pi_{\theta}\right)_{\theta\in[0,1]}$, along which we will apply annealing-based MCMC for sampling from $\pi$. For now, we do not specify the exact form of this curve, but instead introduce the following mild regularity assumption:
@@
Each $\pi_{\theta}$ has a finite second-order moment, and the curve $(\pi_{\theta})_{\theta\in[0,1]}$ is AC with finite action ${\cal A}=\int_{0}^{1}|\dot{\pi}|_{\theta}^{2}\mathrm{d}\theta$.
For the purpose of non-asymptotic analysis, we further introduce the following mild assumption on the target distribution $\pi\propto{\rm e}^{-V}$ on $\mathbb{R}^{d}$, which is widely used in the field of sampling (see Chewi (2024) for an overview):
@@
The potential $V$ is $\beta$-smooth, and there exists a global minimizer $x_{*}$ of $V$ such that $\|x_{*}\|\leq R$. Moreover, $\pi$ has finite second-order moment.
With this foundational setup, we now proceed to introduce the annealed LD and annealed LMC algorithms. Our goal is to characterize the non-asymptotic complexity of annealed MCMC algorithms for sampling from possibly non-log-concave distributions.
@@ 4 Analysis of Annealed Langevin Diffusion
To elucidate the concept of annealing more clearly, we first consider the annealed Langevin diffusion (ALD) algorithm, which samples from the $\pi\propto{\rm e}^{-V}$ by running LD with a dynamically changing target distribution. For the sake of this discussion, we assume the following: (i) we can exactly sample from $\pi_{0}$, (ii) the scores $(\nabla\log\pi_{\theta})_{\theta\in[0,1]}$ are known in closed form, and (iii) we can exactly simulate any SDE with known drift and diffusion terms.
Fix a sufficiently long time $T$. We define a reparametrized curve of probability measures $(\widetilde{\pi}_{t}:=\pi_{t/T})_{t\in[0,T]}$. Starting with an initial sample $X_{0}\sim\pi_{0}=\widetilde{\pi}_{0}$, we run the SDE
$\mathrm{d}X_{t}=\nabla\log\widetilde{\pi}_{t}(X_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},~{}t\in[0,T],$ (3)
@@
When choosing $T=\frac{{\cal A}}{4\varepsilon^{2}}$, it follows that $\operatorname{KL}(\pi\|\nu^{\rm ALD})\leq\varepsilon^{2}$.
@@
Let $\mathbb{Q}$ be the path measure of ALD (Eq. 3) initialized at $X_{0}\sim\widetilde{\pi}_{0}$, and define $\mathbb{P}$ as the path measure corresponding to the following reference SDE:
$\mathrm{d}X_{t}=(\nabla\log\widetilde{\pi}_{t}+v_{t})(X_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},~{}X_{0}\sim\widetilde{\pi}_{0},~{}t\in[0,T].$ (4)
The vector field $v=(v_{t})_{t\in[0,T]}$ is designed such that $X_{t}\sim\widetilde{\pi}_{t}$ for all $t\in[0,T]$. According to the Fokker-Planck equation [^2] Here, we assume that the solution (in terms of the curve of probability measures) of the Fokker-Planck equation given the drift and diffusion terms exists and is unique, which holds under weak regularity conditions (Le Bris & Lions, 2008)., this is equivalent to the following PDE:
@@
which means that $v$ generates $\left(\widetilde{\pi}_{t}:=\pi_{t/T}\right)_{t\in[0,T]}$. We can compute $\operatorname{KL}(\mathbb{P}\|\mathbb{Q})$ using Lem. 1:
$\operatorname{KL}(\mathbb{P}\|\mathbb{Q})=\frac{1}{4}\mathbb{E}_{\mathbb{P}}{\int_{0}^{T}\|v_{t}(X_{t})\|^{2}\mathrm{d}t}=\frac{1}{4}\int_{0}^{T}\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}^{2}\mathrm{d}t.$
Leveraging Lem. 2, among all vector fields $v$ that generate $\left(\widetilde{\pi}_{t}:=\pi_{t/T}\right)_{t\in[0,T]}$, we can choose the one that minimizes $\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}$, thereby making $\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}=|\dot{\widetilde{\pi}}|_{t}$, the metric derivative. With the reparameterization $\widetilde{\pi}_{t}=\pi_{t/T}$, we have the following relation by chain rule:
$|\dot{\widetilde{\pi}}|_{t}=\lim_{\delta\to 0}\frac{W_{2}(\widetilde{\pi}_{t+\delta},\widetilde{\pi}_{t})}{|\delta|}=\lim_{\delta\to 0}\frac{W_{2}(\pi_{(t+\delta)/T},\pi_{t/T})}{T|\delta/T|}=\frac{1}{T}|\dot{\pi}|_{t/T}.$
@@
$\operatorname{KL}(\mathbb{P}\|\mathbb{Q})=\frac{1}{4}\int_{0}^{T}|\dot{\widetilde{\pi}}|_{t}^{2}\mathrm{d}t=\frac{1}{4T}\int_{0}^{1}|\dot{\pi}|_{\theta}^{2}\mathrm{d}\theta=\frac{{\cal A}}{4T}.$
Finally, using data-processing inequality (see, e.g., Chewi (2024, Theorem 1.5.3)), with $T=\frac{{\cal A}}{4\varepsilon^{2}}$,
$\operatorname{KL}(\pi\|\nu^{\rm ALD})=\operatorname{KL}(\mathbb{P}_{T}\|\mathbb{Q}_{T})\leq\operatorname{KL}(\mathbb{P}\|\mathbb{Q})=\varepsilon^{2},$
@@
Let us delve deeper into the mechanics of the ALD. Although at time $t$ the SDE (Eq. 3) targets the distribution $\widetilde{\pi}_{t}$, the distribution of $X_{t}$ does not precisely align with $\widetilde{\pi}_{t}$. Nevertheless, by choosing a sufficiently long time $T$, we actually move on the curve $(\pi_{\theta})_{\theta\in[0,1]}$ sufficiently slowly, thus minimizing the discrepancy between the path measure of $(X_{t})_{t\in[0,T]}$ and the reference curve $(\widetilde{\pi}_{t})_{t\in[0,T]}$. Using the data-processing inequality, we can upper bound the error between the marginal distributions at time $T$ by the joint distributions of the two paths. In essence, moving more slowly, sampling more precisely.
Our analysis primarily addresses the global error across the entire curve of probability measures, rather than focusing solely on the local error at time $T$. This approach is inspired by Dalalyan & Tsybakov (2012); Chen et al. (2023), and stands in contrast to the isoperimetry-based analyses of LD (e.g., Vempala & Wibisono (2019); Chewi et al. (2022); Balasubramanian et al. (2022)), which focus on the decay of the KL divergence from the distribution of $X_{t}$ to the target distribution, and require LSI to bound the time derivative of this quantity. Notably, the total time $T$ needed to run the SDE depends solely on the action of the curve $(\pi_{\theta})_{\theta\in[0,1]}$, obviating the need for assumptions related to log-concavity or isoperimetry.
It is also worth noting that ALD plays a critical role in the field of non-equilibrium stochastic thermodynamics (Seifert, 2012). Recently, a refinement of the fluctuation theorem was discovered in Chen et al. (2020); Fu et al. (2021), showing that the irreversible entropy production in a stochastic thermodynamic system is equal to the ratio of a similar action integral and the duration of the process $T$, closely resembling Thm. 1.